### v1

# Modal Verification Patterns for Systems

Abstract
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources.
In this paper, we argue the position that modalities, in the sense of modal logic, should be a go-to approach when specifying and verifying low-level systems code. We explain how the concept of a resource context helps guide the design of new modalities for verification of systems code, and we justify our perspective by discussing prior systems that have used modalities for systems verification successfully, arguing that they fit into the verification design pattern we articulate, and explaining how this approach might apply to other systems verification challenges.
1 Introduction
Low-level systems exhibit certain patterns in their designs, especially when interacting with computing resources. Exploiting certain patterns while designing software has been an important field of study. In this regard, we think that certain properties of modalities enable us to understand and do the verification challenges that show certain patterns.
1.1 Contributions to Identifying the Verification Patterns via Modal Abstractions
We argue how modal abstractions can be used to identify and abstract system verification challenges. We justify our perspective by discussing prior systems that have successfully used modalities for system verification, arguing that they fit into the verification design pattern we articulate, and explaining how this approach might apply to other systems’ verification challenges.
Identifying System Verification Challenges
We start with identifying common patterns in system verification: virtualization, sharing, and translation.
Introducing the Concept of Resource
Then we discuss the concept of resource which has already been an essential concept in the design of systems. Inspired by the concept of resource in the systems, we define what a resource and its context are in the modal abstractions.
Introducing the Concept of Nominals
Nominalization enables identifying a resource in a context. For example, a transaction is a context of resources of in-memory updated disk blocks. The transaction identifier is used to associate a transaction with a disk-block to be persisted so that, in case of a crash while persisting updated disk-blocks, the filesystem can rollback the already persisted disk-blocks of the transaction, and reach to the previous consistent disk state. To be able to do so, both the updated in-memory disk blocks and the transaction must refer to the transaction identifier – strong nominalization. In another example, virtual memory references (resources) in an address space (resource context), which can be uniquely identified with a root address (the nominal) of its page-table tree, are agnostic to the address space that they are in. However, they can only be accessed (be valid) in the address space to which they are agnostic. However, then an address space switch happens, the virtual memory references of the previous address space must be made inaccessible. To be able to do so, although virtual memory references do not hold any piece of information related to their address space, we still have to associate them. We call this kind of unilateral nominalization – weak nominalization.
Taxonomy of the Current Modal Approaches in System Verification
Based on these concepts defined, we summarize contemporary verification efforts using modal abstractions. We choose them from different domains, for example, weak memory verification, storage persistence  [5, perennialgit, tejperennial19, 27, 8, 9, 6, 19, 28] because we would like to justify that our definitions are not domain dependent.
2 Definitions
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources. In this section, we argue the position that modalities should be a go-to approach when specifying and verifying low-level systems code. We explain how the concept of a resource context helps guide the design of new modalities for verification of systems code, and we justify our perspective by discussing prior systems that have used modalities for systems verification successfully, arguing that they fit into the verification design pattern we articulate, and explaining how this approach might apply to other systems verification challenges.
To explain our ideas in the general systems understanding, we briefly recap some of the background and themes our ideas build on again,casting them in a certain way to bring out the relevance of our philosophy.
2.1 Resources in Systems Software
Systems software, in general, interfaces with an underlying computing architecture such that any software system at any higher level in the software stack can (at least indirectly) utilize the resources of the machine. The last layer of software before the hardware is naturally critical to the correctness of an overall system, as essentially all software built on top of it assumes its correctness. And because hardware is complex and highly diverse, the implementation of those lowest layers of the software stack is typically intricate and naturally error-prone, despite how critical its correctness is. Typically systems software has, as a primary focus, the task of abstracting from hardware details to simplify the construction of higher layers of the stack.
2.1.1 Virtualization
One of the most common forms of abstraction provided by systems software is virtualization, which abstracts the relationship between conceptual and physical computing resources. Operating System (OS) kernels virtualize memory locations and quantity (via virtual memory and paging [7]). Distributed language runtimes may virtualize addresses, even when processes may migrate across machines [17]. Filesystems virtualize locations on disk [24, 16, 25, 4]. Programs built on top of the corresponding systems software layer work logically at the level of these virtualized resources, and it makes sense to specify the systems software directly in terms of those abstractions.
Access by Translation
Accessing virtualized resources via translation is a common way to virtualize notions of location (e.g., virtual memory addresses, inodes or object IDs instead of disk addresses [4, 16, 25, 24]). B-trees , page tables , and related structures both behave like maps, when the corresponding physical resources exist as such just in a different location. Control over the lookup process (e.g, handling the case of a missing translation entry) allows for additional flexibility, such as filling holes in sparse files, or demand paging (both from disk or lazily populating anonymous initially-zeroed mappings). Although the realization of these maps may differ from a system to system based on the context (and sometimes hardware details), they are semantically — logically — partial maps, worth treating as such in verification.
2.2 Nominals
2.2.1 Recapping Modal Operators in Program Specifications: Systems Perspective
The value of modal operators in program specifications is that they permit describing something that is true within some context(s), without needing to fully describe the context itself. Yet unlike shorthands, abbreviations, or general definitions, modal operators interact with the logic in systematic — rather than ad hoc — ways.
For example, most modal operators $M$ have the property that if $P\Rightarrow Q$, then $M(P)\Rightarrow M(Q)$. Many modal operators distribute over conjunction, so $M(P\wedge Q)\Rightarrow(M(P)\wedge M(Q))$. So when a program state or behaviour can be captured by a modality, this interacts nicely with the rest of the logic in ways that simplify both specifications and proofs. Here we give a few examples of how this is true in classic work; the next section argues that this is particularly true for systems software, because the concepts and reasoning that arise in low-level software are naturally captured by modalities.
Although we focus on understanding the modal patterns in the program logics themselves for verifying the client programs (Section 3.1), not in the logical machinery they implement, we think that it sheds light on how the modal operator utilized in the machinery of the program logics shows resemblance and can be lifted to the task of client verification.
Temporal Operators
The best known class of modal logics in the systems community is undoubtedly temporal logics (whether LTL [23], CTL [10], TLA [20], or others), due in part to Lamport’s influence [21] leading to its not-infrequent use in specifying distributed algorithms [22, 1].
2.2.2 Nominalization
Some classes of assertions benefit specifically from naming the explicit conditions where they are true (as opposed to simply requiring them to be true somewhere or everywhere as in the most classic modal operators). This naming generally resembles the satisfaction operator of hybrid logic [3, 2]: @$_{\iota}P$ which evaluates the truth of $P$ at the named (Kripke model) state $\iota$. For this reason we refer to the general idea of naming circumstances explicitly as nominalization, even though the examples we discuss are not necessarily actually hybrid logics.
An example utilization of state naming explicitly on the assertion appears in program logics such as Iris [18], which enables encoding of usage protocols (e.g. state transition system) resembling typestate [26, 11] as specifications. Protocol assertions are annotated with the name of the last (abstract) state at which the protocol is ensured. An example more familiar to the systems community is Halpern et al.’s history of adapting modal logics of knowledge to deal with distributed systems [15, 14, 13]. In most of that line of work, $\mathcal{K}_{a}(P)$ indicates that the node $a$ in the system knows or possesses knowledge of $P$ (for example, a Raft node may “know” a lower bound on the commit index). Alternatively, a modality @$_{i}(P)$ may represent that $P$ is true of/at the specific node $i$ [12] (e.g., that node $i$ has stored a certain piece of data to reliable storage). These permit capturing specific concepts relevant to the correctness (and reasoning about correctness) of a certain class of systems, involving facts about specific named entities in the system.
In each of these cases, the fact that these facts are described using modalities with the core modal property $P\Rightarrow Q$, then $M(P)\Rightarrow M(Q)$ means, for example, that if a process $p$ knows that the commit index is greater than 5 (e.g., $\mathcal{K}_{p}(\mathsf{commitIndex}>5)$) no extra work is required to conclude that the node knows it is greater than 3 (i.e., that $\mathcal{K}_{p}(\mathsf{commitIndex}>3)$), because this follows from standard properties of modal operators as described above. In contrast, if verification instead used a custom assertion $\mathsf{minCommitIndex}(5)$ to represent the former knowledge, one would need to separately provide custom reasoning to conclude $\mathsf{minCommitIndex}(3)$.
3 Contingency Decomposition of a System
3.1 Decomposing a System into its Constituents Contingently
Table 1: Modal Decomposition of Program-Logics.
Resource Context Resource Elements Nominalization Context Steps
Post-Crash Modality  [5, perennialgit, tejperennial19] $\Diamond\;P$ $\ell\mapsto_{n}^{\overline{\gamma}}v$ Strong Crash Recovery
[…]
StackRegion Modality  [27] $\overset{ICut^{n}}{\hookrightarrow}\;P$ $\framebox{n}\;\ell\mapsto v$ Strong Alloc and Return to/from stack
Memory-Fence Modality  [8, 9, 6] $\triangle_{\pi}$ and $\triangledown_{\pi}$ $\ell\mapsto v$ Weak Fence Acquire and Release
Address-Space Modality  [19] [r]P $\ell\mapsto v$ Weak Address-Space Switch
Ref-Count Modality  [28] @ P $\ell_{1}\mapsto v$ Weak Allocating, Dropping and Sharing a Reference
*The StackRegion Modality is an instance of NextGen (called the Independence Modality in [27]).
Lots of existing program logics for system verifications have a common structure, which maps to modalities with a couple extra dimensions of design. We summarize our discussion of these logics in Table 1. We discuss, based on examples, common aspects of how we intuitively think about correctness of systems code in many contexts, articulate those pieces, and call out the commonalities across a range of systems.
3.2 Resource
Consider first the address-space abstraction in an OS kernel. An address-space of a process is a container of virtual addresses referencing data in memory. One would expect to have points-to assertion from separation logic to specify ownership of a memory reference pointing to some data. But that ownership is relative to a specific address space — a specific container. We tend to think directly about what is true in an address space, with the simplest piece being an association between a virtual address and the data it points to. We call the simplest piece, in this and other examples, the resource element:
Definition 1 (Resource Elements)
The simplest atomic facts we want to work with in a particular setting, specific to that setting.
By definition, the resource elements are specific to some limited domain or setting. For example, knowing that a certain address points to a 32-bit signed integer representing 3 is knowledge restricted to a certain address space. In general, we call these domains that any resource element is tied to resource contexts:
Definition 2 (Resource Context)
A resource context is an abstraction, context, or container of resource elements of the same type, e.g., an address space of a process.
We discuss a range of examples for each of these in turn.
Table 1 gives additional examples of systems and their corresponding resource elements and contexts where these elements reside, though none of the work in that table analyzes itself according to the structure we are giving.
Except for Post-Crash-Modality, one can think of the resource contexts in the first column in Table 1 as containers for the corresponding resource elements in the second column.
Stack Regions
When reasoning about stack frame contents, the resource element would be a stack-memory points-to assertion ( $\framebox{n}\;\ell\mapsto v$) indicating that a certain offset into stack region $n$ holds value $v$.
Virtual Memory
For virtual memory management, a virtual-points-to ownership assertion pairing a virtual address ($\ell$) with data ($v$) in an address space is natural. A process’s address space with the root address r is an abstraction that is treated as a container for virtual address mappings, $\ell\mapsto v$;
Weak-Memory
When considering weak memory models, we also want points-to information (address-value mappings).
Reference Counting
When dealing with reference-counting APIs, we may care to specify reachability of memory nodes ($\ell\mapsto v$) in a certain context defined by a shared root address. A shared memory address $\ell$ can be the root of the graph that can be a container of memory nodes ($\ell_{1}\mapsto v$) that are reachable from the root $\ell$.
Post-Crash
The resource element of Post-Crash Modality is not obvious in Table 1, and needs a bit of explanation. Perennial, based on the Iris logic, has both disk-points-to assertions $d[\textsf{p}]\mapsto_{n}v$ (for a specific disk $d$) and in-memory points-to assertions $\ell\mapsto_{n}^{\overline{\gamma}}$. Perennial crash-recovery logic book-keeps resource names (can be thought of as logical variables) $\overline{\gamma}$ to identify which assertions (resource elements) remain valid after a crash — these assertions are only usable while the names in $\overline{\gamma}$ are valid, and a crash resets them, discarding assumptions about volatile state. A subtlety of the notion of a resource context is that, unlike the earlier examples, the context does not need to be a literal data structure. It can instead be (various forms of) a set of executions, as in the Post-Crash and NextGen modalities. The Post-Crash modality $\diamond$ expresses that the assertion $P$ will be true after a crash discards all unstable storage (i.e., RAM). This was the inspiration for the NextGen modality, which is in fact a framework for defining “after-$t$” modalities, where $t$ is an transformation of the global state. [^1] The transformations are subject to some technical constraints that are unrelated to our point here.
3.3 Nominalization
Finally, another design point is the question of whether or not resource element assertions must explicitly track their corresponding context, or if they implicitly pick up their context from where they are used.
Strong nominalization is the case where resource elements must explicitly include the identity of their intended context, while weak nominalization occurs when the resource elements implicitly pick up the relevant context from how they are used. The first three modalities in Table 1 are strongly nominalized, with the resource elements generally carrying identifiers of a specific modality usage. [^2] The Post-Crash modality does not look like this in the presentation here; technically the definition of $\diamond$ quantifies over names $\overline{\gamma}$ internally, dealing with sets of possible contexts. The last three are weakly nominalized.
This choice trades off complexity against flexibility and scoping constraints. Strongly nominalized elements track additional specifier/prover-visible book-keeping data. But in exchange for carrying those identifiers of their context with them, strongly nominalized elements can be used together under any modality. For example, one can use the StackRegion modality to talk about two different stack frames simultaneously for code which accesses multiple stack frames: $\framebox{n}\;\ell\mapsto v\ast\framebox{n+1}\;\ell^{\prime}\mapsto v^{\prime}$. Using a given strongly nominalized assertion element under different modalities for different frames does not change its meaning.
By contrast, weakly nominalized elements are more concise, but then make talking about multiple contexts together marginally more complex: changing which modality an assertion is used with drastically changes its meaning. In the case of the Ref-Count modality, @$_{\ell}(\ell_{1}\mapsto v)$ says that $\ell$ points to a reference count wrapping $\ell_{1}$, while placing the $\ell_{1}\mapsto v$ under a jump modality for a different location entails talking about a different region of memory.
In general, use cases where code frequently manipulates small parts of multiple contexts together should prefer strong nominalization, while use cases where usually larger portions of a single context are at issue should prefer weak nominalization.
4 Conclusion
We have argued the essential patterns that help to choose the modalities when working out how to specify different kinds of systems code based on recent successes. It also captures (a slightly more organized version of) our own thinking in coming up with designs for specifying distributed systems [12], virtual memory managers [19], and in ongoing work using different modalities for different parts of a copy-on-write filesystem (e.g., for assertions true in different snapshots). So not only are the modalities useful for conducting proofs, the pieces we have identified seem to be an effective way of working out a specific modality for a specific use-case.
As our main conclusion and future work, we think that these essential patterns in the designs of modal abstractions discussed constitute the fundamentals of verification patterns when working out how to specify different kinds of systems code.